# Modelling subsonic turbulence with SPH-EXA

### v1

@@
Abstract
The numerical simulation of subsonic turbulence with SPH has traditionally been hindered by E0 errors, inaccurate gradient evaluations, and excessive dissipation. In this work, we present numerical results of SPH simulations of subsonic turbulence and compare to state-of-the-art codes such as AREPO and GIZMO. For the first time, SPH can reproduce the results of such codes, showing a similar interval of wavenumber in the inertial range of the Kolmogorov cascade in the subsonic regime. We use the SPH-EXA code to perform these simulations, a GPU-based state-of-the-art SPH code with high performance and extreme scalability at its core. SPH-EXA is coupled with a modern SPH solver based on methods such as an integral approach to gradient calculation, artificial viscosity switches that include a linear field cleaner, a flexible family of pairing-resistant interpolation kernels, generalized volume elements, and a controlled handling of density jumps which maximizes Lagrangian compatibility. In addition, it includes a novel and extremely scalable gravity solver for astrophysical applications.
I Introduction
Smoothed Particle Hydrodynamics (SPH) is a very flexible and adaptable numerical technique for solving partial differential equations in continuum mechanics that has been successfully applied in many different fields. In particular, its application in fluid dynamics, solid mechanics, and astrophysics has substantially benefited those respective fields since its conception around year 1980 [1, 2]. At the core of its success lay its excellent conservation properties, the lack of an underlying structured mesh, and its natural treatment of three-dimensional calculations [3, 4, 5]. Despite these considerable advantages, SPH has historically struggled with certain types of calculation. In particular, those related to the development of subsonic turbulence.
Turbulence is a fundamental multiscale phenomenon in fluid dynamics, characterized by chaotic and unpredictable fluid flow [6]. It occurs in both natural environments and technological applications, making its study and understanding crucial across various fields of science and engineering. Turbulence plays a vital role in the formation and evolution of astronomical structures, from accretion disks around black holes to the large-scale structure of the universe. It influences the distribution of galaxies, star formation, and the dynamics of the interstellar medium. In meteorology, turbulence affects weather patterns, cloud formation, and the dispersion of pollutants. Understanding turbulence is essential for accurate weather forecasting, climate modeling, and assessing environmental impacts. Ocean currents exhibit turbulent behavior, affecting marine ecosystems, climate systems (through heat and carbon exchange), and navigation. Turbulence also plays a role in phenomena such as El Niño and the mixing of nutrients from deep waters to the surface. In aerospace, automotive, and chemical engineering, controlling or exploiting turbulence can lead to more efficient designs. For example, reducing turbulence around aircraft wings or vehicle bodies can reduce fuel consumption and emissions. In contrast, enhancing turbulence in reactors can improve mixing and chemical reactions. Turbulence challenges our understanding of nonlinear dynamical systems and fluid mechanics. It is a rich field for theoretical and computational research, seeking universal laws and models to describe turbulent flows. Despite its ubiquity and importance, turbulence remains one of the most challenging areas of physics to model and predict due to its complex, non-linear nature. Progress in this field beyond the heuristic but very successful pioneering model by Kolmogorov [7] can lead to advances in technology, improved environmental management, and deeper insights into the fundamental processes of nature.
@@ I Introduction
The numerical simulation of turbulence has been a very powerful tool to progress in its understanding, yet such simulations are very challenging as they require to be accurate across a vast range of scales for very long simulation times. This implies high-resolution simulations with state-of-the-art codes. In general, there are no significant discrepancies between simulations with SPH and with grid-based methods for supersonic turbulence [8]. Nevertheless, reproducing subsonic or transonic turbulence with SPH has been traditionally a challenge because it failed to produce results with sufficient quality to compete with grid-based methods for a similar resolution [9]. First studies of subsonic turbulence with SPH in the context of cosmological simulations [10, 11] which found a rough agreement with the Kolmogorov power spectra but only in a very limited range of wavenumbers. More recent attempts to simulate turbulence by direct numerical simulation (DNS) only led to minor [9] or moderate improvements [12], with the latter highlighting the importance of reducing the viscosity away from shocks. However, the velocity power spectra still showed significant differences with those obtained with mesh-based calculations. All these studies identified zeroth order, $E0$, errors, excessive dissipation, and inaccurate estimation of gradients as the principal causes of the difficulties in simulating turbulence with SPH at low-Mach numbers.
Adapting Large Eddy Simulation (LES) concepts to SPH was shown to be a promising route to simulating turbulence [13, 14] although at the price of introducing additional, and somehow artificial, terms in the basic SPH formulation with a loss in the Lagrangian compatibility of the method. On another note, there is little doubt that DNS simulations of subsonic turbulence would benefit from a better estimation of gradients, [15], e.g. by using an integral approach (IA) to derivatives [16] to obtain a better agreement with the Kolmogorov power spectra and in a wider range of wavelengths.
@@
In this work, we present SPH simulations of subsonic turbulence calculated with the SPH-EXA code, a state-of-the-art, highly scalable hydrodynamic code that incorporates the latest improvements in High-performance Computing (HPC), such as full GPU porting of the SPH computation kernels, efficient domain decomposition, GPU-based neighbor search, on-the-fly initial conditions generation, and in-situ visualization. It also incorporates the latest improvements in the SPH technique such as the mentioned IA to calculate gradients, generalized volume elements, optimized Lagrangian compatibility [17, 18] and an efficient linear field cleaner [19] among others. Equipped with such a bag of resources we show that the simulations of turbulence with SPH-EXA can compete with those obtained with top reference codes such as GIZMO [20] and AREPO [21]. Our results reproduce an increasing inertial range of the Kolmogorov cascade with resolution, which, thanks to the excellent scaling properties of SPH-EXA, will make an increasingly larger fraction of the inertial range of turbulence accessible to DNS in the forthcoming years.
II Methodology
II-A Gradients
@@
SPH-EXA implements an integral approach to calculate derivatives [16, 22].
@@
where the $\tau$ tensor elements are defined as:
$\tau_{ij,a}=\sum_{b}V_{b}~{}(x_{i,b}-x_{i,a})(x_{j,b}-x_{j,a})~{}W_{ab}(h_{a});~{}i,j=1,3\,.$ (2)
@@
$I_{i,a}=\left[\sum_{b}V_{b}f_{b}~{}(x_{i,b}-x_{i,a})~{}W_{ab}(h_{a})\right];~{}i=1,3\,.$ (3)
As a consequence, the gradients calculation is:
$\nabla_{i}f_{a}=\sum_{b}V_{b}(f_{b}-f_{a})A_{i,ab}(h_{a});~{}i=1,3\,$ (4)
@@
$A_{i,ab}(h_{a})=\sum_{j=1}^{3}c_{ij,a}(h_{a})(x_{j,b}-x_{j,a})W_{ab}(h_{a})\,$ (5)
where the summation is up to 3 for 3D simulations and $c_{ij,a}$ the elements of the inverse of matrix $\mathcal{T}$, that is, $\mathcal{C}=\mathcal{T}^{-1}$ (Eq. 1).
II-B Artificial Viscosity
The basic AV implementation in SPH-EXA is the following:
$\Pi_{ab}^{AV}=\begin{cases}-\frac{\alpha}{2}v_{ab}^{sig}~{}w_{ab}&\text{for~{}~{}$\mathbf{x}_{ab}\cdot\mathbf{v}_{ab}<0$}\\ 0&\text{otherwise}\end{cases}\,,$ (6)
where $\alpha$ is the viscous parameter and is usually set to 1. The signal velocity estimator between a pair of particles is:
@@ II-B Artificial Viscosity
A linear velocity field implies a smooth flow without the presence of shocks. Therefore, linear velocity fields should have a vanishing AV. We reconstruct the local velocity field for each particle by removing the linear component of the velocity jump that is used to trigger and calculate the AV in Eq. 6 [19]. Assuming that it is differentiable in the neighborhood of each SPH particle, the best linear approximation of the velocity field is given by multiplying its Jacobian by the distance to its neighbors. Instead of using $\mathbf{v}_{ab}=\mathbf{v}_{a}-\mathbf{v}_{b}$ to calculating $w_{ab}$, we use $\mathbf{v}^{\prime}_{ab}=\mathbf{v}^{\prime}_{a}-\mathbf{v}^{\prime}_{b}$, defined as:
@@
where $\mathbf{x}^{T}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b}$ is the column vector of relative distance and $\mathbf{J}_{\mathbf{v}}$ is the Jacobian of the velocity field. In the case of the presence of discontinuities, such a linear estimator is not accurate enough; hence, the $\phi_{ab}$ factor is designed to go to zero in these cases, while it is kept close to 1 when the field is smooth. To this end, we use a van Leer-like limiter [19]:
@@
$\kappa_{ab}$$\equiv\begin{dcases}\exp{\left[-\left(\frac{q_{ab}-q_{crit}}{q_{fold}}\right)^{2}\right]}&\text{if~{}~{}$q_{ab}<q_{crit}$}\,,\\ 1&\text{if~{}~{}$q_{ab}\geq q_{crit}$}\,,\end{dcases}$ (11)
@@
where $q_{a}$ is the normalized distance between neighbors ($q_{a}=|\mathbf{x}_{b}-\mathbf{x}_{a}|/h_{a}$), $q_{crit}$ is the average of the interparticle distance normalized to the smoothing length in 3D, and $nb_{a}$ is the number of neighbors of particle $a$. The parameter $q_{fold}$ determines the width of the correction factor $\kappa_{ab}$ (Eq. 11) that is applied to the van Leer limiter (Eq. 10) and we use the recommended value [19] of $q_{fold}=0.2$.
The final step is to effectively control excessive dissipation. This can be done by allowing a dynamic individual viscous parameter $\alpha_{a}$ in Eq. 6, rather than having a constant value for all SPH particles. Several methods have been developed to calculate $\alpha_{a}$, but the nominal works of [23] and [24] are the inspiration behind our implementation in SPH-EXA. The main idea is that $\alpha_{a}$ should be initialized by default at a rather low value (usually $\alpha_{a}=0.05$). Then it increases rapidly if dissipation is needed and decays exponentially to its original value when it is not needed, effectively acting as a switch for the AV at the position of each SPH particle.
We start by calculating a local viscous parameter $\alpha_{loc,a}$, following the definition [23]:
$\alpha_{loc,a}=\begin{dcases}\frac{\alpha_{max}\mathcal{D}_{a}}{\mathcal{D}_{a}+h_{a}|\nabla\cdot\mathbf{v}_{a}|+0.05c_{a}}&\text{for~{}~{}$\nabla\cdot\mathbf{v}_{a}<0$}\,,\\ 0&\text{otherwise}\,,\end{dcases}$ (15)
@@
In the case that $\alpha_{a}$ has to decay, we evaluate an instantaneous decay rate $\dot{\alpha}_{a}$ as in [23]:
$\dot{\alpha}_{a}=\begin{dcases}(\alpha_{loc,a}-\alpha_{a})/\tau_{a}&\text{if~{}~{}$\alpha_{min}<\alpha_{loc,a}<\alpha_{a}$}\,,\\ (\alpha_{min}-\alpha_{a})/\tau_{a}&\text{if~{}~{}$\alpha_{min}\geq\alpha_{loc,a}$}\,.\end{dcases}$ (16)
@@
II-C Volume elements
In the context of SPH, where we are discretizing the SPH integral interpolant, it becomes necessary to compute the volume element that corresponds to each particle. An uncomplicated yet highly effective method is $dx^{\prime 3}\equiv V_{b}\sim m_{b}/\rho_{b}$. However, it is known [25, 26] that the discretization of the volume element, present in most SPH equations, can be improved beyond this classical approximation using generalized volume elements (GVE).

### v2

@@
11institutetext: Center for Scientific Computing - sciCORE, University of Basel, Klingelberstrasse 61, 4056 Basel, Switzerland
11email: ruben.cabezon@unibas.ch
22institutetext: Center for Data Analytics - CeDA, University of Basel, Klingelberstrasse 61, 4056 Basel, Switzerland 33institutetext: Departament de Física. Universitat Politècnica de Catalunya (UPC). Av. Eduard Maristany 16, 08019 Barcelona, Spain
33email: domingo.garcia@senior.upc.edu
44institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), 08860 Castelldefels (Barcelona), Spain 55institutetext: Universität Heidelberg, Zentrum für Astronomie, Institut für Theoretische Astrophysik, Albert-Ueberle-Str. 2, 69120 Heidelberg, Germany 66institutetext: Department of Mathematics and Computer Science, University of Basel, Spiegelgasse 1, 4051 Basel, Switzerland 77institutetext: Swiss National Supercomputing Center (CSCS), Via Trevano 131, 6900 Lugano, Switzerland 88institutetext: Department of Astrophysics, University of Zürich, Winterthurerstrasse 190, 8057 Zürich, Switzerland 99institutetext: Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Im Neuenheimer Feld 225, 69120 Heidelberg, Germany
From particles to precision
Simulating subsonic turbulence with Smoothed Particle Hydrodynamics
Rubén M. Cabezón    Domingo García-Senz    Oliver Avril    Osman Seckin Simsek    Sebastian Keller    Axel S. Lechuga    Lucio Mayer    Ralf Klessen    Florina M. Ciorba
(Received September 30, 20XX)
Direct numerical simulation of subsonic turbulence with smoothed particle hydrodynamics (SPH) has traditionally been hampered by zeroth-order (E0) errors, inaccurate gradient evaluations, and excessive numerical dissipation. We demonstrate that the recently developed SPH-EXA code can overcome these challenges, producing results similar to those of other state-of-the-art non-SPH codes, such as AREPO and GIZMO. We show that SPH-EXA accurately reproduces the Kolmogorov inertial range scaling in the subsonic regime with increasing resolution, closely matching state-of-the-art non-SPH methods. This is an outstanding result given the important role of SPH codes in astrophysics and cosmology, and the ubiquity of turbulent phenomena.
Key Words.:
hydrodynamics – turbulence – Methods: numerical
[…]
1 Introduction
Smoothed Particle Hydrodynamics (SPH) is a mesh-free, conservative method for continuum mechanics widely used in many different fields since its conception around 1980 (Lucy, 1977; Gingold and Monaghan, 1977). Its strengths include excellent conservation properties and a natural treatment of fully three-dimensional flows (Liu and Liu, 2010; Price, 2012b). Despite these advantages, SPH has historically struggled in some regimes, most notably in accurately simulating subsonic turbulence.
Turbulence is a fundamental multiscale phenomenon in fluid dynamics, characterized by chaotic and nonlinear motion (Frisch, 1995). It occurs in both natural environments and technological applications, making its study central across science and engineering. In astrophysics, turbulence influences processes ranging from accretion flows to the interstellar medium, shaping star formation, galaxy evolution, and the emergence of large-scale structure. Despite its ubiquity and importance, turbulence remains one of the most challenging phenomena to model and predict because of its complex, non-linear nature. Progress in this field beyond the heuristic but very successful pioneering model of Kolmogorov (Kolmogorov, 1941) can lead to advances in technology, improved environmental management, and deeper insight into fundamental processes of nature.
@@ 1 Introduction
The direct numerical simulation (DNS) of turbulence requires high-resolution and state-of-the-art codes. In general, there are no significant discrepancies between SPH simulations and grid-based methods for supersonic turbulence (Kitsionas et al., 2009). Nevertheless, reproducing subsonic or transonic turbulence with SPH has traditionally been a challenge because it failed to produce results with sufficient quality to compete with grid-based methods for a similar resolution (Bauer and Springel, 2012). The first studies of subsonic turbulence with SPH in the context of cosmological simulations (Dolag et al., 2005; Valdarnini, 2011) found rough agreement with the Kolmogorov power spectra but only in a very limited range of wavenumbers. More recent DNS attempts led to minor (Bauer and Springel, 2012) or moderate improvements (Price, 2012a), with the latter highlighting the importance of reducing the viscosity away from shocks after showing that the Reynolds number is proportional to the Mach number in SPH simulations. However, the velocity power spectra still showed significant differences with those obtained with mesh-based calculations. All of these studies identified zeroth-order errors ($E0$), excessive dissipation, and inaccurate gradient estimation as the main causes of the difficulties in simulating turbulence with SPH in low-Mach numbers.
In this work, we present SPH simulations of subsonic turbulence with SPH-EXA [^1] https://github.com/sphexa-org/sphexa
https://deepwiki.com/sphexa-org/sphexa
@@
(Cavelan et al., 2020; Keller et al., 2023), a GPU-native state-of-the-art and highly scalable hydrodynamic code that incorporates the latest advances in high performance Computing (HPC), such as efficient domain decomposition, GPU-based neighbor search, on-the-fly initial conditions generation, and in-situ visualization. It also incorporates the latest improvements in the SPH technique, such as an integral approach to calculate gradients, generalized volume elements, and a slope-limited reconstruction (SLR) of the velocity field to control dissipation, among other improvements (see Appendices A and B). Equipped with such a set of resources, we show that the simulations of subsonic turbulence with SPH-EXA can compete with those obtained with top reference codes such as AREPO (Springel, 2010) and GIZMO (Hopkins, 2015). Our results show that, as resolution increases, we recover a growing portion of the inertial range of the Kolmogorov cascade. Owing to the excellent scaling properties of SPH-EXA, this will allow DNS to probe an ever larger fraction of the turbulent inertial range in the coming years.
Figure 1: Comparison of the power spectra of the velocity field with different elements incorporated into the SPH solver. The dashed line is the theoretical Kolmogorov power spectrum of $\sim k^{-5/3}$. STD denotes standard SPH and is taken from (Bauer and Springel, 2012, simulation S3 in their Fig. 6, with $256^{3}$ particles). All other simulations were performed with SPH-EXA and $250^{3}$ particles. The meaning of the remaining acronyms is provided in Sect. 3.
Figure 2: Comparison of the instantaneous power spectra of the velocity field at $t=10$ at increasing resolution. Each line corresponds to an increase of a factor of 2 in resolution with respect to the previous curve.
2 Initial conditions
We have designed an embedded initial conditions (IC) generator in SPH-EXA. In all our tests, we use a template of a relaxed box with $50^{3}$ particles. This box has previously been relaxed to achieve a glass-like configuration (Arth et al., 2019) with constant density and periodic boundary conditions. Replicating this box in different directions allows us to create on-the-fly initial conditions with a user-defined number of particles and in a distributed manner, only limited by the memory of the system.
All simulations use a periodic box of side $L=1$ with uniform initial density $\rho=1$ and pressure $P=1$, and ’quasi’-isothermal equation of state ($\gamma=1.001$). Turbulence is driven by stochastic forcing modes at different wavelengths k in the range $1<|\mathbf{k}|<3$, with amplitudes following a parabolic distribution $\propto 1-(k-k_{c})^{2}$ peaking at $k_{c}=2$, ensuring that kinetic energy is injected only at the largest scales of the spectrum. The resulting forcing is built from a mixture of $1/3$ compressive and $2/3$ solenoidal modes, whose phases evolve following a stochastic Ornstein-Uhlenbeck process (Uhlenbeck and Ornstein, 1930; Federrath et al., 2010). Simulations start from homogenous ICs and are driven for $t=10$ (in units of sound-crossing time, $t_{sc}=L/c=1$), until a statistically stationary state with RMS Mach number $\mathcal{M}\sim 0.34$ is reached.
3 Results
We present the main results of the simulations carried out with SPH-EXA. Technical details, such as the integral approach to derivatives (IA) used to evaluate gradients, the use of generalized volume elements (GVE), the treatment of dissipation with both variable AV coefficients (VAV), and slope reconstruction technique (SLR), are deferred to the appendices. The appendices also include additional diagnostics based on the second-order velocity structure function, highlighting the advantages of SPH’s fully Lagrangian description.
3.1 Impact of solver elements at fixed resolution
To evaluate the weight and significance of the different elements that contribute to our SPH implementation, we conducted identical simulations while progressively incorporating more of these elements. Figure 1 shows the power spectrum of the velocity field of simulations of $250^{3}$ particles with increasing complexity, compared to the S3 simulation of (Bauer and Springel, 2012) with $256^{3}$ particles and the classic SPH formulation (i.e. gradients calculated directly as kernel derivatives, no GVEs, and no viscosity switches), labeled STD in our plot.
Figure 3: Slice of thickness $2h$ centered on $z=0$ showing the velocity magnitude with increasing resolution (SPH-EXA). We represent particles directly.
The inclusion of IA recovers the correct slope in the inertial range of the Kolmogorov cascade at low $k$, considerably extending it compared to STD (Valdarnini, 2016), but it does not remove the overly steep slope $\sim-4$ at large $k$. At these small scales something else is needed. Adding GVE and especially VAV moves in the right direction, revealing the bottleneck effect as an energy excess near the resolution limit, as commonly seen in mesh codes. Finally, the slope-limited reconstruction (SLR) provides the decisive improvement: it extends the inertial range almost a decade relative to STD, makes the bottleneck clearly visible, and yields a dissipative-range slope comparable to non-Lagrangian codes (see Fig. 5).
3.2 Resolution scaling and flow morphology
In Fig. 2 we present the power spectrum of the velocity field for simulations performed with SPH-EXA and with different particle counts at the end of each simulation ($t=10$). It is clear that with increasing resolution, the inertial range over the Kolmogorov scale is extended to larger wavenumber, and more importantly, that the rate of enhancement with increasing number of particles is not marginal or small, but considerable and consistent with that obtained with modern mesh-based and moving-mesh methods. This is a novel result, which means that SPH can indeed resolve smaller scales with higher resolution, in opposition to the results found with older SPH implementations (e.g., Bauer and Springel, 2012). Figure 3 shows a thin slice of the velocity field for increasing resolution with SPH-EXA, from $200^{3}$ up to $3200^{3}$, where it is evident how we can resolve a delicate and rich mixture of large- and small-scale features. Figure 4 compares the mass-weighted and volume-weighted density PDFs, with a fitted normal distribution. We can see the expected tail bias towards low densities, since in subsonic flows, local pressure increases in colliding flows can prevent the buildup of high-density regions (Konstandin et al., 2012).
Figure 4: Instantaneous density PDF for the $400^{3}$ simulation at $t=10$. Comparison of mass-weighted and volume-weighted PDFs.
3.3 Cross-code comparison
We can now compare our results directly to other state-of-the-art codes: moving-mesh and meshless finite-volume methods. Figure 5 shows velocity power spectra from AREPO, GIZMO, and SPH-EXA. All three reproduce a comparable inertial range and a bottleneck of similar magnitude at comparable $k$. This suggests that, once the injection method and effective viscosity are matched, the cascade between the forcing and the dissipation wavenumbers behaves similarly across methods. Note that we compare runs with a similar number of fluid elements, which does not necessarily imply the same spatial resolution. In SPH, resolution is set by the smoothing length, which covers significantly more than just one particle. At $256^{3}$ cells, a proxy is the cell size, which for AREPO is $\sim 1/256\sim 3.9\times 10^{-3}$, comparable to our $400^{3}$ SPH simulation, with an average of $h\sim 3.6\times 10^{-3}$. Spatial resolution matters the most in the dissipation regime, where AREPO, GIZMO, and SPH-EXA have different transfer functions. In particular, GIZMO and SPH-EXA show a similar gentler slope in the roll-off, while AREPO is slightly steeper, pointing to a stronger high-$k$ damping. However, these differences are modest and the qualitative behavior persists at higher resolution, with the break shifting to larger $k$, as expected.
Figure 5: Comparison of the velocity power spectra for different codes at $t=10$. AREPO and GIZMO curves are from their respective papers.
Figure 6 shows, from top to bottom, the module of the velocity, density, and enstrophy of the fluid at the end of our simulations. The two columns on the left are an excerpt of the results with AREPO and the SPH used in (Bauer and Springel, 2012). It is clear from these two columns that the SPH used in that work was unable to resolve subsonic turbulence because it lacks all small-scale features. The column on the right shows the results of SPH-EXA with $250^{3}$ particles. The similarities with the results of AREPO are evident, and there is a rich mix of fine structures. This can also be seen in Fig. 3, which illustrates a slice of the module of the velocity field at $t=10$ for different resolutions. SPH-EXA also produces an excellent second-order velocity structure function (appendix C). Additional colormaps of the highest resolution simulation, $3200^{3}$, can be found in appendix D.
Figure 6: Comparison of SPH-EXA results with AREPO and SPH-tav in Bauer and Springel (2012) at $t=25$. From top to bottom: module of velocity, density, and enstrophy. The two columns on the left are an excerpt of the original results of AREPO and their SPH with reduced artificial viscosity (both with $256^{3}$ fluid elements) of Bauer and Springel (2012) shown in their Fig. 4. The column on the right is the result of SPH-EXA with $250^{3}$ particles. The fields show directly particles in a thin slice around $Z=0$. No rasterization was performed.
4 Summary and conclusions
SPH-EXA combines a modern SPH formulation with state-of-the-art parallelization and GPU acceleration. This enables simulations that mitigate historical SPH limitations, such as excessive dissipation and inaccurate gradient evaluations, that have long hindered SPH from matching the performance of grid-based and moving-mesh codes in this field. Our findings not only align well with results from non-SPH codes like AREPO and GIZMO, but also illustrate the robustness of SPH in handling complex turbulent flows at resolutions previously thought unattainable with this method.
We demonstrate that, with increasing resolution, the inertial range of the power spectra extends to larger wavenumbers, exhibiting features comparable to those seen in non-SPH codes. These improvements allow SPH to better exploit its intrinsic strengths: truly Lagrangian sampling (particles are the fluid elements, avoiding tracer advection and particle-to-grid interpolation noise), excellent conservation properties (exact mass advection and pairwise-conservative exchanges of momentum and energy), and flexibility across physical regimes. In particular, SPH couples naturally and accurately to tree methods for self-gravity, supporting its use in more demanding astrophysical contexts.
Acknowledgements.
This work has been supported by the Swiss Platform for Advanced Scientific Computing (PASC) project ”SPH-EXA: Optimizing Smoothed Particle Hydrodynamics for Exascale Computing”. It has also been carried out as part of the SKACH consortium through funding from SERI. We also acknowledge the financial support of the European Research Council through the ERC Synergy Grant “ECOGAL” (project ID 855130), from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) “STRUCTURES”, and from the German Ministry for Economic Affairs and Climate Action in project “MAINN” (funding ID 50OO2206). The numerical calculations have been supported by the EuroHPC-JU Extreme Scale Access Mode through the project ”TGSF: The role of turbulence and gravity in star formation” (EHPC-EXT-2023E01-031). The authors acknowledge the support of the LUMI Supercomputer Center, the Swiss National Supercomputing Center (CSCS - allocations c32 and ch18), and the Center for Scientific Computing (sciCORE) at the University of Basel.
Appendix A SPH Ingredients
In this section, we discuss the ingredients of our SPH formulation that have the largest impact on our simulations of subsonic turbulence. Namely, the calculation of gradients, dissipation, and generalized volume elements.
A.1 Gradients
@@
SPH-EXA implements an integral approach to calculate derivatives (García-Senz et al. 2012; Cabezón et al. 2012).
@@
where the $\tau$ tensor elements are defined as the weighted sum of the dyadic product of the relative distances $(\mathbf{x}_{b}-\mathbf{x}_{a})\otimes(\mathbf{x}_{b}-\mathbf{x}_{a})$:
$\tau_{ij,a}=\sum_{b}V_{b}\penalty 10000\ (x_{i,b}-x_{i,a})(x_{j,b}-x_{j,a})\penalty 10000\ W_{ab}(h_{a});\penalty 10000\ i,j=1,3\,.$ (2)
@@
$I_{i,a}=\left[\sum_{b}V_{b}f_{b}\penalty 10000\ (x_{i,b}-x_{i,a})\penalty 10000\ W_{ab}(h_{a})\right];\penalty 10000\ i=1,3\,.$ (3)
As a consequence, the gradient calculation is:
$\nabla_{i}f_{a}=\sum_{b}V_{b}(f_{b}-f_{a})A_{i,ab}(h_{a});\penalty 10000\ i=1,3\,$, (4)
@@
$A_{i,ab}(h_{a})=\sum_{j=1}^{3}c_{ij,a}(h_{a})(x_{j,b}-x_{j,a})W_{ab}(h_{a})\,$, (5)
where the summation is up to 3 for 3D simulations and $c_{ij,a}$ the elements of the inverse of tensor $\mathcal{T}$, that is, $\mathcal{C}=\mathcal{T}^{-1}$ (Eq. 1).
A.2 Dissipation
The basic dissipation control in SPH-EXA is done using the following artificial viscosity (AV) prescription:
$\Pi_{ab}^{AV}=\begin{cases}-\frac{\alpha}{2}v_{ab}^{sig}\penalty 10000\ w_{ab}&\text{for\penalty 10000\ \penalty 10000\ $\mathbf{x}_{ab}\cdot\mathbf{v}_{ab}<0$}\\ 0&\text{otherwise}\end{cases}\,$, (6)
where $\alpha$ is the viscous parameter and is set to 1. The signal velocity estimator between a pair of particles is:
@@ A.2 Dissipation
A linear velocity field implies a smooth flow without the presence of shocks. Therefore, linear velocity fields should have a vanishing AV. We reconstruct the local velocity field for each particle by removing the linear component of the velocity jump that is used to trigger and calculate the AV in Eq. 6 (Frontiere et al. 2017). This is the slope-limited reconstruction (SLR) method. Assuming that it is differentiable in the neighborhood of each SPH particle, the best linear approximation of the velocity field is given by multiplying its Jacobian by the distance to its neighbors. Instead of using $\mathbf{v}_{ab}=\mathbf{v}_{a}-\mathbf{v}_{b}$ to calculating $w_{ab}$, we use $\mathbf{v}^{\prime}_{ab}=\mathbf{v}^{\prime}_{a}-\mathbf{v}^{\prime}_{b}$, defined as:
@@
where $\mathbf{x}^{T}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b}$ is the column vector of relative distance and $\mathbf{J}_{\mathbf{v}}$ is the Jacobian of the velocity field. In the case of discontinuities, such a linear estimator is not accurate enough; hence, the $\phi_{ab}$ factor is designed to go to zero in these cases, while it is kept close to 1 when the field is smooth. To this end, we use a van Leer-like limiter (Frontiere et al. 2017):
@@
[…]
@@
where $q_{a}$ is the normalized distance between neighbors ($q_{a}=|\mathbf{x}_{b}-\mathbf{x}_{a}|/h_{a}$), $q_{crit}$ is the average of the interparticle distance normalized to the smoothing length in 3D, and $nb_{a}$ is the number of neighbors of particle $a$. The parameter $q_{fold}$ determines the width of the correction factor $\kappa_{ab}$ (Eq. 11) that is applied to the van Leer limiter (Eq. 10) and we use the recommended value (Frontiere et al. 2017) of $q_{fold}=0.2$.
The final step is to control excessive dissipation effectively. This can be done by allowing a dynamic individual viscous parameter $\alpha_{a}$ in Eq. 6, rather than having a constant value for all SPH particles. Several methods have been developed to calculate $\alpha_{a}$, but the nominal works of Cullen and Dehnen (2010) and Read and Hayfield (2012) are the inspiration behind our implementation in SPH-EXA. The main idea is that $\alpha_{a}$ should be initialized by default at a rather low value (usually $\alpha_{a}=0.05$). Then it increases rapidly if dissipation is needed and decays exponentially to its original value when it is not needed, effectively acting as a switch for the AV at the position of each SPH particle.
We start by calculating a local viscous parameter $\alpha_{loc,a}$, following the definition (Read and Hayfield 2012):
[…]
@@
In the case that $\alpha_{a}$ has to decay, we evaluate an instantaneous decay rate $\dot{\alpha}_{a}$ as in (Read and Hayfield 2012):
$\dot{\alpha}_{a}=\begin{dcases}(\alpha_{loc,a}-\alpha_{a})/\tau_{a}&\text{if\penalty 10000\ \penalty 10000\ $\alpha_{min}<\alpha_{loc,a}<\alpha_{a}$}\,,\\ (\alpha_{min}-\alpha_{a})/\tau_{a}&\text{if\penalty 10000\ \penalty 10000\ $\alpha_{min}\geq\alpha_{loc,a}$}\,.\end{dcases}$ (16)
@@
In this respect, it is worth noticing the latest works pointing to the interesting possibility of not needing switches at all, if a modulated version of the SLR is used (Price 2024; Sandnes et al. 2025; García-Senz and Cabezón 2025). This is achieved by taking constant $\alpha=1$ and multiplying the limiter $\phi_{ab}$ in Eqs. 8 and 9 by:
$1-\mathcal{B}^{p}\,$, (17)
where $\mathcal{B}$ is the Balsara limiter,
${\mathcal{B}}_{a}=\frac{|\nabla\cdot\mathbf{v}|_{a}}{|\nabla\cdot\mathbf{v}|_{a}+|\nabla\times\mathbf{v}|_{a}+10^{-4}\frac{c_{a}}{h_{a}}}\,.$ (18)
García-Senz and Cabezón (2025) showed that taking $p=2$ in Eq. 17 provides the best results across a series of tests involving shocks and/or shear, improving considerably over the use of switches and switches+SLR techniques. However, in order to get the best results in simulating subsonic turbulence, they recommend including an additional control in the amplitude of the AV to avoid spurious triggering of dissipation by the random motion of particles in the smaller scales. This can be done either with a switch, as shown before, or directly with a limiter, such as Balsara.
A.3 Volume elements
In the context of SPH, where we are discretizing the SPH integral interpolant, it becomes necessary to compute the volume element that corresponds to each particle. An uncomplicated yet highly effective method is $dx^{\prime 3}\equiv V_{b}\sim m_{b}/\rho_{b}$. However, it is known (Saitoh and Makino 2013; Hopkins 2013) that the discretization of the volume element, present in most SPH equations, can be improved beyond this classical approximation using generalized volume elements (GVE).